Brian asked a group of people their favourite holiday destination.
The results are summarised in the table.

Destination UK Europe USA Africa Other
Frequency 84 144 108 24 120
Angle

Work out the size of each angle to draw a pie chart.
1
Expert's answer
2020-11-29T19:10:48-0500

Total number of people

"84+144+108+24+120=480"


UK:

"\frac{84}{480}\cdot360\degree=63\degree"


Europe:

"\frac{144}{480}\cdot360\degree=108\degree"


USA:

"\frac{108}{480}\cdot360\degree=81\degree"


Africa:

"\frac{24}{480}\cdot360\degree=18\degree"


Other:

"\frac{120}{480}\cdot360\degree=90\degree"